OpenStudy (anonymous):

sorry this is a lot to ask for. but I need to study for my final! If someone could help me out it would be awesome! this is on sequence and series!

OpenStudy (anonymous):

PLEASE HELP ME

OpenStudy (anonymous):

1.) what are the first five terms of the sequence given by the formula An=5n+1? A) 1,5,9,13,17 B) 6,11,16,21,26 C) 5,10,15,20,25 D) 5,6,7,8,9 2)write a recrusive definition for the sequence 11,8,5,2,... A) a1=11,an=an-1-4 B) a1=11,an=an-1-3 C) a1=2, an=an-1-3 D) a1=2, an=an-1+3 3) write an explicit definition for the sequence -3,1,5,9,... A) an=-3+4(n-1) B) an=-3+4n C) an=-3+ (n-1) D) an= -3+4 (n+1) 4) suppose you drop a tennis ball from a height of 8 feet. after the ball hits the floor, it rebounds to 80% of its previous height. how high will the the ball rebound after its 3rd bounce? A) 3.3 ft B) 4.1 ft C 5.1 ft D) 1 ft 6) is the sequence arithmetic? if so identify the common difference. 14,21,42,77,... A) yes; 7 B) yes;-7 C) yes; 14 D) no 7) what is the 12th term of the sequence 15,11,7,... A) -29 B) -25 C)-21 D) -33 8) find the missing term of the arithmetic sequence 24, ___ , 36,... A)34 B)30 C)42 D)20 10) what is the 10th term in the sequence 625,125,25,... A) 1/3125 B) 1/625 C) 1/1,250 D) 1/15,625 11) what is a possible value for the missing term of the geometric sequence? 39, __ , 975,... A)-44 B)-195 C)-5 D)-4,875 12) what is the sum of the finite arithmetic series? (-10) + 0+ 20+..... + 130 A) 900 B) 1,000 C) 950 D) 955 14) what is the sum of the arithmetic series? \[\sum_{t=1} (3t-4)\] A)234 B)441 C)25.5 D)49 15) what is the sum of the first 10 terms of the sequence defined by an=3n-3? A)132 B)138 C)27 D)135 16) what is the sum of the geometric series rounded to the nearest whole number? \[\sum_{x=0}^{15} 2(1/3)^x\] A)2 B)3 C)4 D)1 18) does the infinite geometric series diverge or converge? Explain A) diverge; it does not have a sum B) converges; it does not have a sum C) diverges; it has a sum D) converges; it has a sum

OpenStudy (anonymous):

Mai Gawd O_O

OpenStudy (noseboy908):

Let's start with number one, and we'll see how far we can get. So, number one states "1.) what are the first five terms of the sequence given by the formula An=5n+1? A) 1,5,9,13,17 B) 6,11,16,21,26 C) 5,10,15,20,25 D) 5,6,7,8,9 "

OpenStudy (noseboy908):

Do you know how to find the terms, given n?

OpenStudy (anonymous):

Well kinda. not quite! these are just all my question off my practice test I don't understand!

OpenStudy (noseboy908):

Ok, then think of each one like this: If n is equal to "x", then just put in whatever value x is for the "n" variable in "5n+1". Example: If n=1 (As your first one is), then put in one for n, getting "5(1)+1", which is the same as 6. Understand?

OpenStudy (anonymous):

yea, that makes sence. do you always use 1 in replacement for x?

OpenStudy (noseboy908):

No, you replace n with whichever term you're trying to find. If you want to find the third term, for example, then just put a 3 in place of n.

OpenStudy (anonymous):

oh! okay. I got it!

OpenStudy (noseboy908):

Then what's the answer for one?

OpenStudy (anonymous):

B

OpenStudy (noseboy908):

Correct! Now, number two!

OpenStudy (noseboy908):

2)write a recrusive definition for the sequence 11,8,5,2,... A) a1=11,an=an-1-4 B) a1=11,an=an-1-3 C) a1=2, an=an-1-3 D) a1=2, an=an-1+3

OpenStudy (anonymous):

now, on this one can you help me get started I struggle on these.

OpenStudy (noseboy908):

I'm just thinking, one moment.

OpenStudy (anonymous):

okay, thank you for even helping me! take your time.

OpenStudy (noseboy908):

Ok, firstly, what's the first term listed?

OpenStudy (anonymous):

its 11

OpenStudy (noseboy908):

Yep, so that narrows the possible answers down to two, as saying "a1" is simply the first term. So, given this, it must be which of two answers? What or what?

OpenStudy (anonymous):

its either A or B?

OpenStudy (noseboy908):

Yep. That being said, what's being added or subtracted from one term to get the next one?

OpenStudy (anonymous):

its gotta be B?

OpenStudy (noseboy908):

Correct, as three is being subtracted, not four. Time for number three!

OpenStudy (noseboy908):

3) write an explicit definition for the sequence -3,1,5,9,... A) an=-3+4(n-1) B) an=-3+4n C) an=-3+ (n-1) D) an= -3+4 (n+1)

OpenStudy (anonymous):

awesome! sounds good

OpenStudy (noseboy908):

The root formula for arithmetic equations will help greatly here. It is "a(n)=a1+d(n-1)", where a1 is the first term, d is how much is added or subtracted from a term to get the next one, and where n is the nth term in a series. So, we can ignore the a1 part, because all the answers have the same beginning parts. Ok?

OpenStudy (anonymous):

ok, lets try this!

OpenStudy (noseboy908):

Now, what do we know is being added or subtracted from one term to the next?

OpenStudy (anonymous):

adding 4

OpenStudy (noseboy908):

Yep! So which variable I just listed does that fit the description of?

OpenStudy (anonymous):

would it be d?

OpenStudy (noseboy908):

You're close, but look at the last part of it. There's one difference between it and the root formula. What is it?

OpenStudy (anonymous):

sorry! its A, I just noticed it.

OpenStudy (noseboy908):

There ya go :), correct! Make sure the sign is the same as the root formula before deciding the answer. Time for number 4!

OpenStudy (noseboy908):

4) suppose you drop a tennis ball from a height of 8 feet. after the ball hits the floor, it rebounds to 80% of its previous height. how high will the the ball rebound after its 3rd bounce? A) 3.3 ft B) 4.1 ft C 5.1 ft D) 1 ft

OpenStudy (anonymous):

okay I will! perfect.

OpenStudy (noseboy908):

Essentially, we're finding 80%(80%(80% of 8 (feet)))). Understand where I got this?

OpenStudy (anonymous):

yea. I do!

OpenStudy (noseboy908):

Good, then tell me, what is 80% of 80% of 80% of 8?

OpenStudy (anonymous):

80% of 8 is 6?

OpenStudy (noseboy908):

No, how'd you get that? Make sure that you're doing ".8 x 8"

OpenStudy (anonymous):

is the answer D?

OpenStudy (noseboy908):

Nope.

OpenStudy (noseboy908):

What is 80% of 8?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

hm, I not sure how this one works

OpenStudy (noseboy908):

Put this into a calcuator: .8 x 8. What is the result?

OpenStudy (anonymous):

.64

OpenStudy (noseboy908):

Yep. Now, multiply that by .8 again.

OpenStudy (anonymous):

0.512

OpenStudy (noseboy908):

Oh, go back and do .8 x 8 again please. Make sure you note the placement of the decimal.

OpenStudy (anonymous):

6.4?

OpenStudy (noseboy908):

Yeah, that's it. Now, multiply that by .8.

OpenStudy (anonymous):

5.12

OpenStudy (noseboy908):

And once more please!

OpenStudy (anonymous):

the answer is B!

OpenStudy (noseboy908):

Correct!

OpenStudy (noseboy908):

Now, problem 5! 6) is the sequence arithmetic? if so identify the common difference. 14,21,42,77,... A) yes; 7 B) yes;-7 C) yes; 14 D) no

OpenStudy (noseboy908):

There is no number 5, so number 6 instead!

OpenStudy (anonymous):

? @noseboy908

OpenStudy (anonymous):

yea, just straight to number 6 hah

OpenStudy (noseboy908):

Now, for number 6, the key is that Arithmetic equations will always have the same "d". Thus, each term must be found by adding the same number each time to find the next one. (In all, if one number minus the previous number does NOT equal any other given number minus the previous number, then it's not arithmetic). So, what is added to the first term to get the second?

OpenStudy (anonymous):

7 is added

OpenStudy (anonymous):

A?

OpenStudy (noseboy908):

Don't be too hasty, let's check another two terms. Please tell me what is being added to the second term to get the second term.

OpenStudy (anonymous):

21

OpenStudy (noseboy908):

Mhmm. Is that the same as what's being added to a1 to get a2?

OpenStudy (anonymous):

no its not

OpenStudy (noseboy908):

Yep, so what do we then know about the series? Doesn't an arithmetic sequence require a constant number for "d"?

OpenStudy (anonymous):

correct. its not an arithmetic sequence

OpenStudy (noseboy908):

There ya go. Time for number 7!

OpenStudy (anonymous):

lol yup!

OpenStudy (noseboy908):

7) what is the 12th term of the sequence 15,11,7,... A) -29 B) -25 C)-21 D) -33

OpenStudy (noseboy908):

Firstly, please restate the root formula for arithmetic equations that I previously gave you.

OpenStudy (anonymous):

one sec. I think I can do this one, just let me know if im wrong.

OpenStudy (noseboy908):

Ok.

OpenStudy (anonymous):

the answer would be A ?

OpenStudy (noseboy908):

Yep! At that, I'm afraid I must go, at least for now. But that's correct. I should return later to help if I'm still needed.

OpenStudy (anonymous):

dang! around what time will you be back? I need your help! or the answers and I will work till I figure them out.

OpenStudy (noseboy908):

It's likely that I'll be back before 7 CST, ok?

OpenStudy (anonymous):

so like 4 hours? I will message you then?

OpenStudy (noseboy908):

No, about two hours.

OpenStudy (anonymous):

oh, okay. well then I will message you in 2 hours! thank you!

OpenStudy (noseboy908):

8) find the missing term of the arithmetic sequence 24, ___ , 36,... A)34 B)30 C)42 D)20

OpenStudy (noseboy908):

Are you ready for more?

OpenStudy (anonymous):

hey, im sorry! im back @noseboy908

OpenStudy (noseboy908):

Alright, then let's begin!

OpenStudy (anonymous):

sweet. im ready, sorry I just sent you a messege

OpenStudy (noseboy908):

To start, please restate the root formula I gave you that is for arithmetic equations.

OpenStudy (anonymous):

have the same "d". Thus, each term must be found by adding the same number each time to find the next one. (In all, if one number minus the previous number does NOT equal any other given number minus the previous number, then it's not arithmetic). is it this one?

OpenStudy (noseboy908):

Look again, I'm talking about a specific formula. It starts with "a(n)=a1..."

OpenStudy (anonymous):

"a(n)=a1+d(n-1)"

OpenStudy (noseboy908):

Exactly! So, plug in what you know. What are you given, in terms of variables?

OpenStudy (anonymous):

with 24 being the first variable?

OpenStudy (noseboy908):

Be specific: which variable?

OpenStudy (anonymous):

n?

OpenStudy (anonymous):

the answer is b correct?